2.)Solve the system using substitution, explain your steps
3x – 4y = 10
2y = 2x – 6

Answer by josgarithmetic(39618)

Solve one of the equations for either variable; substitute the expression for this variable into the other equation. Solve for the variable in this equation. Use the result or the value to solve for the other variable.

Example, find x from your first equation:
3x-4y=10
3x=10+4y
x=10/3+4y/3
-
Substitute into second equation:
2y=2(10/3+4y/3)-6
y=(10/3+4y/3)-3
3y=10+4y-1, after dividing by three
0=10-1+y
-9=y
-
Use the value of y to find value of x:
x=10/3+4(-9)/3
x=(10-36)/3
x=-26/3
-
-
ANSWER: x=-26/3 and y=-9